• Main INDEX
  • Monthly INDEX
  • PREV
    Make New Entry, Make Followup Entry

    User name Ramesh/Diancheng

    Log entry time 21:58:52 on October 31, 2009

    Entry number 297075

    keyword=Left HRS DAQ work

    Besides explicitly rechecking T2, all triggers are working the way they
    should and are also plugged in the normalization scaler. There are two
    LeCroy scalers before the normalization scaler. After the normalization
    scaler also there two scalers, the first of which has gas-cherenkov
    signals in the ribbon cable in scaler ch0-15 and S1 signals in the
    ribbon cable in scaler ch16-31, the second scaler has s2m left and
    right PMT signls in its 32 channels.

    ADC gate was set to be 300 ns wide so as to put the pion rejector
    signal within the gate, whereas the TDC gate is kept only as 200 ns.
    Some randomly chosen S1, S2m and pion rejector ADC signals were checked
    and seen within the ADC gate.

    Gas-cherenkov PMT #7 has low rate (~200 Hz) compared to other PMT like
    #5 (400 Hz) and #9 (400 Hz).

    S2m paddle left PMT #10 to #15 have the same type of signal with the
    same rate, hence, we were not able to see any noisy PMT there (as Vince
    was suggesting looking at the histogram, problem may be elsewhere).

    In the PVDIS rack, veto signal is 50 ns earlier than the T1 signal. We
    were not able to see a working veto today (plan to continue working
    tomorrow morning). A cosmic run is underway overnight though.